Suppose x is a uniform random variable with a = 1 and b = 27. Find the mean. 35 O 70 3.35 O 14
Answer #1

Answer: 14

>> For a uniform random variable with the parameters (a, b), the value of the mean is (a + b)/2. Where, a is the minimum value and b is the maximum value of the random variable.

Thus, the mean is (a + b)/2 = (1 + 27)/2 = 14.